12 pts./True or false section(Write down BOTH QUESTION ANDANSWER at your answer booklet DONT GUESS You throw a dice and toss a coin simultaneously. The sample space of this experiment contains 8 different values. a. t 12 b. ACRV has uniform distribution from-4 to 8. Its variance is 4/3. c. | The probability of a certain event is 1. d. If A and B are two disjoint events,where PIA1-0.1 and PIB-0.4 then PIA or BI-0.5 e. The event A and its complement A are independent events f A fair die is rolled six times. The probability that the highest outcomeof the six throws (T) 1p | (T) I is less than 4 is 1/6. A CRV has uniform distribution from 5 to 3. The probability that the CRV is positive is 3/8 A jar contains 3 red marbles, 7 green marbles and 10 white marbles. If a marble is drawn from the larat random, the probability that this marble is white 0.5. ()Ip h. iThe slope of the PDF for CRV couldnt be negative. r Ip . Ir A and B are two independent events, where PIA)-0.2 and P[B)-0.5 then PIA, B-0.7 (F) Ip CE IP k. Two mutually-exclusive events are independent. L The number of different 8-bit words containing three ones and five zeros are 40.

With explanation please.
Answer #1

a)

dice - 6 different possibility

coin - 2 different possibility

hence sample space = 6*2 = 12

FALSE

b)

variance = (b-a)^2/12

= (8 +4)^2/12 = 12

FALSE

c)

An event is said to be a Certain Event if there is an absolute surety on the chance of its occurrence. The probability of certain event is said to be 1.

TRUE

d)

when A and B are disjoint

P(A and B) = 0

P(A or B) = P(A) + P(B) - P(A and B)

= P(A) + P(B)

= 0.1 + 0.4

=0.5

TRUE

Please rate

Please post rest questions again